# A Comprehensive Survey for Hyperspectral Image Classification: The Evolution from Conventional to Transformers

### v3

@@ 2 Hyperspectral Data Representation
HS data is typically a 3D hypercube, denoted as $\textbf{X}\in\mathcal{R}^{B\times(N\times M)}$, containing spectral and spatial information of a sample [57]. Here, $B$ is the number of spectral bands, and $N$ and $M$ represent the spatial components (width and height) [58]. Figure 2 illustrates an HSI cube from the University of Houston dataset.
@@
In spectral representation, each pixel vector, denoted as $x_{i}\in\mathcal{R}^{B}$, is processed independently based on its spectral signature. The value of $B$ represents the spectral channels, which can be the actual number or a reduced set obtained through dimensionality reduction (DR) techniques. A lower-dimensional representation is preferred to reduce redundancy and enhance class separability without significant information loss [59]. Unsupervised DR techniques, like Principal Component Analysis (PCA) [60], and Locally Linear Embedding [61], transform HSI data without using class labels. Supervised DR methods, such as linear discriminant analysis (LDA) [62], local Fisher discriminant analysis (LFDA) [63], local discriminant embedding (LDE) [64, 65, 66], and nonparametric weighted feature extraction (NWFE) [67], utilize labeled samples to improve class separability. LDA and LFDA, for instance, enhance class separability by maximizing the distance between classes while minimizing within-class distance. However, spectral mixing challenges differentiation between classes solely based on spectral reflectance. This underscores the need for advanced techniques beyond spectral information for accurate HSI class differentiation.
@@
To overcome spectral representation limitations, spatial information of pixels can be utilized. In this approach, pixels in each spectral band are represented as a matrix, denoted as $x_{i}\in\mathcal{R}^{N\times M}$. Spatial representation considers neighboring pixels due to high spatial correlation, enhancing class coherence. Neighbors are identified using kernels or pixel-centric windows [68]. Various methods extract spatial information from the HSI cube, including morphological profiles (MPs) [59], Gabor filters for texture features [69], gray-level co-occurrence matrix (GLCM) [70], local binary patterns (LBP) [71], and Deep Neural Network (DNN)-based methods. MPs capture geometrical characteristics, with extensions like extended morphological profiles (EMPs) [59] and multiple-structure-element morphological profiles [72]. Texture extraction methods like Gabor filters capture textural details, while LBP [73] provides rotation-invariant representations. GLCM [70, 74] determines spatial variability based on relative pixel positions. DNNs can extract spatial information by treating each pixel as an image patch, learning spatial features. Combining multiple methods, as seen in studies like Zhang et al. [75], enhances spatial information extraction. In this study, Gabor filters and differential MPs were combined for an RNN-based HSC framework, showcasing the potential of integrating techniques for leveraging spatial information.
@@
Spatial-spectral approaches in HSI process pixel vectors by incorporating both spectral features and spatial-contextual information. Two common strategies exist for the simultaneous use of spectral and spatial representations. The first strategy concatenates spatial details with the spectral vector, demonstrated in studies by Chen et al. [76, 77], merging both into a single feature representation. The second strategy processes the three-dimensional HSI cube, preserving its structure and contextual information. Paoletti et al. [78] showcase this approach, maintaining the HSI’s three-dimensional nature and utilizing DL techniques to extract features capturing both spectral and spatial characteristics. Both strategies aim to enhance HSI analysis and classification tasks by leveraging the complementary nature of spectral and spatial information. While many DNN models focus on spectral representation [79], recognizing its limitations, recent efforts aim to integrate spatial information into the classification process [79]. Joint exploitation of spectral and spatial features has gained interest, leading to improved accuracy [80].
@@ 3 Learning Categories
Supervised Learning In supervised learning, models are trained on labeled data, associating each example with a target output. The model adjusts its parameters iteratively during training to improve accuracy in predicting outputs. Effective training requires a substantial amount of labeled data, making these models suitable for scenarios with abundant labeled data. DNNs excel in learning complex patterns and making accurate predictions in such scenarios. Unsupervised Learning Unsupervised learning involves training models on unlabeled data to uncover patterns without predefined target labels. Measuring accuracy can be challenging, but it is useful when learning intrinsic structures of datasets with limited labeled training data. Techniques like PCA learn low-dimensional representations by identifying directions of maximum variance. K-means clustering groups data into clusters, revealing patterns or similarities among instances [81].
Semi-Supervised Learning strategy combines elements of both supervised and unsupervised learning. It utilizes a limited amount of labeled data along with a larger set of unlabeled data for training. The model learns from the labeled data to make predictions on both labeled and unlabeled instances, leveraging the additional unlabeled data for improved performance [82, 83]. Reinforcement Learning, an agent interacts with an environment and learns to take actions to maximize cumulative rewards. The agent receives feedback in the form of rewards or penalties based on its actions, allowing it to learn optimal decision-making policies through trial and error [84]. Transfer Learning involves leveraging knowledge or representations learned from one task or domain and applying them to a different but related task or domain. This approach can be beneficial when labeled data is scarce or when pre-trained models can provide useful feature representations [85, 86]. Online Learning, also known as, incremental learning, involves updating the model continuously as new data becomes available. The model learns from each instance of data sequentially, adapting to changing patterns or concepts over time [87, 88].
@@
Self-learning for HSC leverages the model’s ability to iteratively improve its performance without extensive labeled data. This approach utilizes techniques such as self-supervised learning, where the model learns from the inherent structure and properties of the HS data. By generating pseudo-labels and employing consistency regularization, the model refines its understanding of spectral-spatial patterns. Self-learning is particularly beneficial in HSI due to the high dimensionality and limited availability of labeled samples, enabling more efficient and accurate classification by maximizing the utility of available data [69, 89, 90, 91, 92].
@@ 4 Brief Overview of Conventional Methods
The main goal of HSC is to assign distinct labels to pixel vectors in the HSI cube based on their spectral or spatial characteristics [93]. Mathematically represented as $\textbf{X}=[x_{1},x_{2},x_{3},\dots,x_{B}]^{T}\in\mathcal{R}^{B\times(N\times M)}$, where $B$ is the total number of spectral bands with $N\times M$ samples per band [94]. These samples belong to classes Y, each $x_{i}=[x_{1,i},~{}x_{2,i},~{}x_{3,i},\dots,x_{B,i}]^{T}$ represents the $i^{th}$ sample with a class label $y_{i}\in\mathcal{R}^{Y}$. The classification task is treated as an optimization problem, where a mapping function $f_{c}(.)$ operates on the input data X [48], aiming to minimize the disparity between the obtained output and actual labels [95] as shown in equation 1.
@@
where the parameter $\theta$ adjusts transformations applied to the input data X for the mapping function $f_{c}:\textbf{X}\to\textbf{Y}$. HSC research has seen a trend influenced by computer vision methodologies [1]. TML-based HSC relies on handcrafted features like shape, texture, color, spectral, and spatial details, and classification techniques. Common feature extraction methods include texture descriptors (LBPs [29], HOG [96], PHOG [97]), SIFT [98], GIST [98] and classifiers include Random Forests [99], kernel-based SVM [100], KNN [101], and ELM [55].
Color histograms are simple and effective but lack spatial context, making them sensitive to illumination changes. HOG captures edge orientations effectively [102]. SIFT is robust but computationally intensive. GIST provides a global image description based on statistical properties like roughness and openness [98]. Texture descriptors like LBPs are common for capturing texture around each pixel [29, 103]. Color histograms, GIST, and texture descriptors are global features, capturing statistical characteristics, while HOG and SIFT are local features describing geometric information [104]. Combining these features is a common practice in HSC [105]. LBPs are computationally efficient, involving simple binary operations on local pixel neighborhoods. In contrast, SIFT is more complex, requiring convolutions, gradient computations, and orientation assignments, making it computationally intensive [106].
Hand-crafted features, while effective, face challenges in real-world data due to varying optimal feature sets, subjectivity in human-driven design, biases, and limitations. Traditional HSC approaches encounter the curse of dimensionality, struggle with feature selection and extraction, lack spatial information consideration, exhibit limited robustness to noise, face scalability issues, and may not adapt well to complex data distributions. These challenges, including the curse of dimensionality, reliance on domain expertise, and limited spatial information, have prompted the exploration of DL-based methods, such as CNNs and RNNs, for HSC. DL techniques offer automated feature learning, potentially overcoming the limitations of traditional methods and improving classification accuracy by capturing intricate spatial-spectral relationships in HSIs [107, 108].
@@ 5 Advantages of Deep Learning
Deep Learning (DL) architectures, exemplified in Figure 4, can discern patterns and features from data without prior knowledge of its statistical distribution [109]. Unlike conventional methods, DL models extract both linear and non-linear features without pre-specified information, making them suitable for handling HS data in spectral and spatial domains, individually or combined [110, 1]. With flexible architectures and varying depths, DL models accommodate different layers, allowing adaptation to diverse machine learning strategies—supervised, semi-supervised, and unsupervised. This adaptability makes DL architectures optimized and tailored for specific classification tasks across various domains and applications [111].
@@
DL enhances HSI data processing, leading to heightened classification accuracy and the ability to capture intricate relationships. DL models automatically learn and extract informative features from high-dimensional HSI data, addressing inherent complexity and variability [112]. Overall, DL architectures present a promising approach for HSC, learning from data without prior distribution knowledge, and offering architectural and learning strategy flexibility [113]. These capabilities make DL models well-suited for overcoming challenges and complexities in HSI data analysis.
@@
Automatic feature learning: DL models autonomously learn relevant features from data, creating hierarchical representations at various abstraction levels [114]. Unlike traditional methods relying on hand-crafted features, DL adapts and extracts discriminative features from HSIs, reducing the need for manual engineering and potentially capturing complex spectral-spatial patterns. Spatial spectral fusion: DL models effectively capture spatial and spectral characteristics by integrating both types of information [115]. CNNs exploit spatial structure through convolutional layers, discerning local patterns [116]. Combining spectral and spatial details enhances accuracy, especially for classes with similar spectra but distinct spatial distributions.
@@
End-to-end learning: DL models are trained end-to-end, optimizing the entire model, including feature extraction and classification, jointly [117, 118]. This simultaneous learning of features and classifiers can lead to superior performance compared to separate stages in traditional approaches. Adaptability to complex data distributions: DL models, particularly CNNs, with their non-linear transformations, effectively capture complex data distributions and intricate relationships between features and classes [119]. This adaptability enables better handling of variability and nonlinearity in HSIs.
@@
Transfer learning and pre-training: DL models benefit from transfer learning, leveraging knowledge from one task or dataset for another [120]. Pre-training on large-scale datasets, followed by fine-tuning on smaller HS datasets, enhances classification performance, especially with limited labeled data [50]. Scalability: DL models efficiently handle large-scale datasets, enabling quick classification for real-time applications. Advances in hardware and DL frameworks, like Graphics Processing Units (GPUs), accelerate training and inference processes, making DL applicable to larger and more complex HS datasets [121, 122]. Despite these advantages, DL models for HSC face challenges, including the need for large labeled datasets, computational requirements, and potential overfitting. However, ongoing developments in DL techniques and increased HS dataset availability mitigate these challenges, establishing DL as a promising approach for accurate and automated HSC.
@@ 6 Convolutional Neural Networks for HSC
The CNN architecture, inspired by Hubel and Wiesel’s biological visual system [123] and Fukushima’s Neocognitron model [124], comprises two main stages: Feature Extraction (FE) and classification. The FE network, with convolutional and pooling layers, extracts hierarchical representations of input data. The subsequent classification stage utilizes these features, demonstrating CNN’s success in tasks like image classification and object detection. The FE network includes stacked convolutional layers, activation, dropout, and pooling—each contributing to meaningful FE.
The convolutional layer, sharing kernels across the input, captures local patterns efficiently, reducing model complexity and enhancing training ease. In a CNN, convolved results undergo non-linear transformations through an activation layer, extracting crucial non-linear features for capturing complex data patterns. Following activation, pooling reduces feature map resolution, achieving shift-invariance and retaining essential information. Typically, a pooling layer follows each convolutional layer, coupled with activation. This hierarchical process extracts increasingly complex features from input data for subsequent analysis and classification. The classification stage incorporates Fully Connected (FC) layers and a Softmax operator to determine class probabilities. Global average pooling, an FC layer alternative, summarizes spatial information and reduces parameters, mitigating overfitting. Softmax normalizes outputs into a class probability distribution, aiding predictions. While Softmax is common, studies explore SVM as an alternative for classification in CNNs [125], offering different optimization objectives and decision boundaries. Overall, the classification stage utilizes FC layers or global average pooling, followed by a Softmax operator or alternatives like SVM, for predicting class memberships in input patterns. Subsequent sections detail three CNN architectures tailored for HSC: i) Spectral CNN, ii) Spatial CNN, and iii) Spectral-spatial CNN.
6.1 Spectral CNN
Spectral CNNs exclusively process 1D spectral data $x_{i}\in\mathcal{R}^{B}$, where $B$ denotes the original or reduced spectral bands [126]. Tailored for HS data, they treat each pixel’s spectral information as a 1D vector, capturing spectral intensity at different wavelengths. Unlike conventional 2D CNNs for image classification, these models consider HSIs as 1D sequences, where each element represents spectral intensity. The input is the pixel’s spectral profile, either the original full-dimensional vector or a dimensionality-reduced representation, presenting a focused approach to handle spectral dimensionality.
@@ 6.1 Spectral CNN
A crucial element in spectral CNNs is the convolutional layer, applying filters to capture local spectral patterns and generate feature maps by element-wise operations. Commonly, pooling layers, such as max or average pooling, downsample and reduce dimensionality. Additional layers like fully connected or recurrent layers can be added for enhanced discriminative power. In [127], a CNN addressed overfitting by using $1\times 1$ convolutional kernels, increased dropout rates, and a global average pooling layer instead of fully connected layers. For handling high correlation among HSI bands, Gao, et al. [128] introduced a CNN architecture transforming the 1D spectral vector into a 2D feature matrix with cascading $1\times 1$ and $3\times 3$ convolutional layers. This design allowed feature reuse across different spectral bands. Similarly, Li et al. [129] employed a global average pooling layer to reduce trainable parameters and extract high-dimensional features.
Wu et al. [130] proposed a hybrid model combining convolutional and recurrent layers for extracting both position-invariant middle-level features and spectral-contextual details. Similarly, Jin et al. [131] proposed a hybrid architecture that effectively classified healthy and diseased wheat heads by transforming spectral information into a 2D structure for improved analysis. In spectral-based identification of rice seed varieties, Qiu et al. [132] demonstrated CNN’s superiority over SVM and KNN, showcasing its effectiveness in accurately identifying rice seed varieties based on spectral characteristics. A similar application in [133] identified various Chrysanthemum varieties using CNN on spectral data represented by the first five Principal Components (PCs) obtained through PCA, a widely-used dimensionality reduction method. In medical HSI, Huang et al. [134] used PCA as a preprocessing step, fusing CNN kernels with Gabor kernels for classification, highlighting the efficacy of hybrid architectures and the superiority of CNN over traditional algorithms in spectral-based classification tasks for HSC.
Advantages: Firstly, Spectral CNN captures local spectral patterns effectively through convolutional operations, facilitating discriminative feature extraction [126]. Secondly, by integrating pooling operations, spectral CNNs manage high-dimensional HS data, reducing computational complexity. Thirdly, the inclusion of fully connected or recurrent layers enables learning high-level representations and capturing spectral-contextual dependencies, enhancing classification accuracy.

### v4

@@ 2 Hyperspectral Data Representation
HS data is typically a 3D hypercube, denoted as $\textbf{X}\in\mathcal{R}^{B\times(N\times M)}$, containing spectral and spatial information of a sample (Ahmad,, 2023). Here, $B$ is the number of spectral bands, and $N$ and $M$ represent the spatial components (width and height). Figure 2 illustrates an HSI cube from the University of Houston dataset.
@@
In spectral representation, each pixel vector, denoted as $x_{i}\in\mathcal{R}^{B}$, is processed independently based on its spectral signature. The value of $B$ represents the spectral channels, which can be the actual number or a reduced set obtained through dimensionality reduction (DR) techniques. A lower-dimensional representation is preferred to reduce redundancy and enhance class separability without significant information loss (Zhang et al., 2024e, ). Unsupervised DR techniques, like Principal Component Analysis (PCA) (Mantripragada and Qureshi,, 2024), and Locally Linear Embedding (Hong et al., 2020c, ), transform HSI data without using class labels. Supervised DR methods, such as linear discriminant analysis (LDA), local Fisher discriminant analysis (LFDA), local discriminant embedding (LDE), and nonparametric weighted feature extraction (NWFE) (Kuo et al.,, 2009), utilize labeled samples to improve class separability. LDA and LFDA, for instance, enhance class separability by maximizing the distance between classes while minimizing within-class distance. However, spectral mixing challenges differentiation between classes solely based on spectral reflectance. This underscores the need for advanced techniques beyond spectral information for accurate HSI class differentiation.
@@
To overcome spectral representation limitations, spatial information of pixels can be utilized. In this approach, pixels in each spectral band are represented as a matrix, denoted as $x_{i}\in\mathcal{R}^{N\times M}$. Spatial representation considers neighboring pixels due to high spatial correlation, enhancing class coherence. Neighbors are identified using kernels or pixel-centric windows (Chhapariya et al.,, 2024). Various methods extract spatial information from the HSI cube, including morphological profiles (MPs) (Zhang et al., 2024e, ), Gabor filters for texture features (Cheng et al.,, 2024), gray-level co-occurrence matrix (GLCM) (Tian et al., 2024b, ), local binary patterns (LBP) (Hong et al.,, 2015), and Deep Neural Network (DNN)-based methods. MPs capture geometrical characteristics, with extensions like extended morphological profiles (EMPs) (Zhang et al., 2024e, ) and multiple-structure-element morphological profiles (Hong et al., 2020a, ). Texture extraction methods like Gabor filters capture textural details, while LBP (Zhang et al., 2024f, ) provides rotation-invariant representations. GLCM (Tian et al., 2024b, ; Tsai and Lai,, 2013) determines spatial variability based on relative pixel positions. DNNs can extract spatial information by treating each pixel as an image patch, learning spatial features. Combining multiple methods, as seen in studies like Zhang et al. (Zhang et al., 2018b, ), enhances spatial information extraction. In this study, Gabor filters and differential MPs were combined for an RNN-based HSC framework, showcasing the potential of integrating techniques for leveraging spatial information.
@@
Spatial-spectral approaches in HSI process pixel vectors by incorporating both spectral features and spatial-contextual information. Two common strategies exist for the simultaneous use of spectral and spatial representations. The first strategy concatenates spatial details with the spectral vector, demonstrated in studies by Chen et al. (Chen et al.,, 2014, 2015), merging both into a single feature representation. The second strategy processes the three-dimensional HSI cube, preserving its structure and contextual information. Paoletti et al. (Paoletti et al., 2018b, ) showcase this approach, maintaining the HSI’s three-dimensional nature and utilizing DL techniques to extract features capturing both spectral and spatial characteristics. Both strategies aim to enhance HSI analysis and classification tasks by leveraging the complementary nature of spectral and spatial information. While many DNN models focus on spectral representation (Chen and Wang,, 2016), recognizing its limitations, recent efforts aim to integrate spatial information into the classification process (Chen and Wang,, 2016). Joint exploitation of spectral and spatial features has gained interest, leading to improved accuracy (Ghamisi et al.,, 2015).
@@ 3 Learning Categories
In supervised learning, models are trained on labeled data, associating each example with a target output. The model adjusts its parameters iteratively during training to improve accuracy in predicting outputs. Effective training requires a substantial amount of labeled data, making these models suitable for scenarios with abundant labeled data. DNNs excel in learning complex patterns and making accurate predictions in such scenarios. Unsupervised Learning Unsupervised learning involves training models on unlabeled data to uncover patterns without predefined target labels. Measuring accuracy can be challenging, but it is useful when learning intrinsic structures of datasets with limited labeled training data. Techniques like PCA learn low-dimensional representations by identifying directions of maximum variance. K-means clustering groups data into clusters, revealing patterns or similarities among instances (Sun et al.,, 2022).
3.b Unsupervised Learning
Unsupervised learning in HSI is a powerful approach for analyzing high-dimensional data without the need for labeled training samples. This method leverages the rich spectral information inherent in HSIs to identify intrinsic patterns and structures within the data. Techniques such as clustering, dimensionality reduction, and feature extraction are commonly employed to uncover underlying data distributions and group similar spectral signatures. For instance, algorithms like k-means clustering or hierarchical clustering can be used to segment regions of interest based on their spectral characteristics, facilitating tasks such as anomaly detection and material identification. Additionally, advanced methods like self-organizing maps, autoencoders, and generative models are increasingly being explored to enhance the interpretability of HSI data, enabling more robust insights into land cover classification and environmental monitoring applications. The ability to extract meaningful information from unlabeled data makes unsupervised learning an essential tool in the HSI domain, particularly in scenarios where obtaining labeled datasets is challenging or costly (Cao et al., 2024b, ; Khan et al.,, 2021).
3.c Semi-Supervised Learning
Semi-supervised learning combines elements of both supervised and unsupervised learning. It utilizes a limited amount of labeled data along with a larger set of unlabeled data for training. The model learns from the labeled data to make predictions on both labeled and unlabeled instances, leveraging the additional unlabeled data for improved performance (Wang et al., 2024e, ; Dópido et al.,, 2013).
3.d Reinforcement Learning
An agent interacts with an environment and learns to take actions to maximize cumulative rewards. The agent receives feedback in the form of rewards or penalties based on its actions, allowing it to learn optimal decision-making policies through trial and error.
3.e Transfer Learning
Transfer learning involves leveraging knowledge or representations learned from one task or domain and applying them to a different but related task or domain. This approach can be beneficial when labeled data is scarce or when pre-trained models can provide useful feature representations (Zhong et al.,, 2020; He et al.,, 2020).
3.f Online Learning
Online learning also known as, incremental learning, involves updating the model continuously as new data becomes available. The model learns from each instance of data sequentially, adapting to changing patterns or concepts over time (Fu et al.,, 2018; Hong et al.,, 2017).
3.g Self-learning
@@
Self-learning for HSC leverages the model’s ability to iteratively improve its performance without extensive labeled data. This approach utilizes techniques such as self-supervised learning, where the model learns from the inherent structure and properties of the HS data. By generating pseudo-labels and employing consistency regularization, the model refines its understanding of spectral-spatial patterns. Self-learning is particularly beneficial in HSI due to the high dimensionality and limited availability of labeled samples, enabling more efficient and accurate classification by maximizing the utility of available data (Cheng et al.,, 2024; Zhao et al., 2024d, ; Li et al., 2024a, ).
@@ 4 Brief Overview of Conventional Methods
The main goal of HSC is to assign distinct labels to pixel vectors in the HSI cube based on their spectral or spatial characteristics (Ahmad et al., 2022b, ). Mathematically represented as $\textbf{X}=[x_{1},x_{2},x_{3},\dots,x_{B}]^{T}\in\mathcal{R}^{B\times(N\times M)}$, where $B$ is the total number of spectral bands with $N\times M$ samples per band Ahmad et al., 2019b . These samples belong to classes Y, each $x_{i}=[x_{1,i},~{}x_{2,i},~{}x_{3,i},\dots,x_{B,i}]^{T}$ represents the $i^{th}$ sample with a class label $y_{i}\in\mathcal{R}^{Y}$. The classification task is treated as an optimization problem, where a mapping function $f_{c}(.)$ operates on the input data X (Ahmad et al., 2021b, ), aiming to minimize the disparity between the obtained output and actual labels (Paoletti et al.,, 2019) as shown in equation 1.
@@
where the parameter $\theta$ adjusts transformations applied to the input data X for the mapping function $f_{c}:\textbf{X}\to\textbf{Y}$. HSC research has seen a trend influenced by computer vision methodologies (Ahmad et al., 2021c, ). TML-based HSC relies on handcrafted features like shape, texture, color, spectral, and spatial details, and classification techniques. Common feature extraction methods include texture descriptors (LBPs, HOG (Anahara,, 2015), PHOG (Ge et al.,, 2023)), SIFT (Al-khafaji et al.,, 2018), GIST (Al-khafaji et al.,, 2018) and classifiers include Random Forests (Zhang et al., 2018c, ), kernel-based SVM (Zhong et al., 2018a, ), KNN (Ma et al.,, 2010), and ELM (Ahmad et al., 2019c, ).
Color histograms are simple and effective but lack spatial context, making them sensitive to illumination changes. HOG captures edge orientations effectively (Hong et al.,, 2016). SIFT is robust but computationally intensive. GIST provides a global image description based on statistical properties like roughness and openness (Al-khafaji et al.,, 2018). Texture descriptors like LBPs are common for capturing texture around each pixel (Wang et al.,, 2023). Color histograms, GIST, and texture descriptors are global features, capturing statistical characteristics, while HOG and SIFT are local features describing geometric information (Gong et al.,, 2018). Combining these features is a common practice in HSC (Hong et al., 2019b, ). LBPs are computationally efficient, involving simple binary operations on local pixel neighborhoods. In contrast, SIFT is more complex, requiring convolutions, gradient computations, and orientation assignments, making it computationally intensive (Nhat and Hoang,, 2019).
Hand-crafted features, while effective, face challenges in real-world data due to varying optimal feature sets, subjectivity in human-driven design, biases, and limitations. Traditional HSC approaches encounter the curse of dimensionality, struggle with feature selection and extraction, lack spatial information consideration, exhibit limited robustness to noise, face scalability issues, and may not adapt well to complex data distributions. These challenges, including the curse of dimensionality, reliance on domain expertise, and limited spatial information, have prompted the exploration of DL-based methods, such as CNNs and RNNs, for HSC. DL techniques offer automated feature learning, potentially overcoming the limitations of traditional methods and improving classification accuracy by capturing intricate spatial-spectral relationships in HSIs (Hinton and Salakhutdinov,, 2006; Hu et al.,, 2015).
@@ 5 Advantages of Deep Learning
Deep Learning (DL) architectures, exemplified in Figure 4, can discern patterns and features from data without prior knowledge of its statistical distribution (Li et al., 2024b, ). Unlike conventional methods, DL models extract both linear and non-linear features without pre-specified information, making them suitable for handling HS data in spectral and spatial domains, individually or combined (Shrestha and Mahmood,, 2019; Ahmad et al., 2021c, ). With flexible architectures and varying depths, DL models accommodate different layers, allowing adaptation to diverse machine learning strategies—supervised, semi-supervised, and unsupervised. This adaptability makes DL architectures optimized and tailored for specific classification tasks across various domains and applications (Venkateswara et al.,, 2017).
@@
DL enhances HSI data processing, leading to heightened classification accuracy and the ability to capture intricate relationships. DL models automatically learn and extract informative features from high-dimensional HSI data, addressing inherent complexity and variability (Chen and Wang,, 2014). Overall, DL architectures present a promising approach for HSC, learning from data without prior distribution knowledge, and offering architectural and learning strategy flexibility (Ahmad et al., 2021a, ). These capabilities make DL models well-suited for overcoming challenges and complexities in HSI data analysis.
@@
Automatic feature learning: DL models autonomously learn relevant features from data, creating hierarchical representations at various abstraction levels (Niu et al.,, 2019). Unlike traditional methods relying on hand-crafted features, DL adapts and extracts discriminative features from HSIs, reducing the need for manual engineering and potentially capturing complex spectral-spatial patterns. Spatial spectral fusion: DL models effectively capture spatial and spectral characteristics by integrating both types of information (Zhong et al., 2021a, ). CNNs exploit spatial structure through convolutional layers, discerning local patterns (Yu et al.,, 2020). Combining spectral and spatial details enhances accuracy, especially for classes with similar spectra but distinct spatial distributions.
@@
End-to-end learning: DL models are trained end-to-end, optimizing the entire model, including feature extraction and classification, jointly (Pan et al.,, 2020; Li et al., 2023c, ). This simultaneous learning of features and classifiers can lead to superior performance compared to separate stages in traditional approaches. Adaptability to complex data distributions: DL models, particularly CNNs, with their non-linear transformations, effectively capture complex data distributions and intricate relationships between features and classes (Liu et al., 2020b, ). This adaptability enables better handling of variability and nonlinearity in HSIs.
@@
Transfer learning and pre-training: DL models benefit from transfer learning, leveraging knowledge from one task or dataset for another (Pan et al.,, 2011). Pre-training on large-scale datasets, followed by fine-tuning on smaller HS datasets, enhances classification performance, especially with limited labeled data (Ahmad et al., 2022a, ). Scalability: DL models efficiently handle large-scale datasets, enabling quick classification for real-time applications. Advances in hardware and DL frameworks, like Graphics Processing Units (GPUs), accelerate training and inference processes, making DL applicable to larger and more complex HS datasets (Ullah et al.,, 2024; Singh et al.,, 2023). Despite these advantages, DL models for HSC face challenges, including the need for large labeled datasets, computational requirements, and potential overfitting. However, ongoing developments in DL techniques and increased HS dataset availability mitigate these challenges, establishing DL as a promising approach for accurate and automated HSC.
@@ 6 Convolutional Neural Networks for HSC
The CNN architecture, inspired by Hubel and Wiesel’s biological visual system (Hubel and Wiesel,, 1962) and Fukushima’s Neocognitron model (Fukushima,, 1980), comprises two main stages: Feature Extraction (FE) and classification. The FE network, with convolutional and pooling layers, extracts hierarchical representations of input data. The subsequent classification stage utilizes these features, demonstrating CNN’s success in tasks like image classification and object detection. The FE network includes stacked convolutional layers, activation, dropout, and pooling—each contributing to meaningful FE.
The convolutional layer, sharing kernels across the input, captures local patterns efficiently, reducing model complexity and enhancing training ease. In a CNN, convolved results undergo non-linear transformations through an activation layer, extracting crucial non-linear features for capturing complex data patterns. Following activation, pooling reduces feature map resolution, achieving shift-invariance and retaining essential information. Typically, a pooling layer follows each convolutional layer, coupled with activation. This hierarchical process extracts increasingly complex features from input data for subsequent analysis and classification. The classification stage incorporates Fully Connected (FC) layers and a Softmax operator to determine class probabilities. Global average pooling, an FC layer alternative, summarizes spatial information and reduces parameters, mitigating overfitting. Softmax normalizes outputs into a class probability distribution, aiding predictions. While Softmax is common, studies explore SVM as an alternative for classification in CNNs (Hong et al.,, 2017), offering different optimization objectives and decision boundaries. Overall, the classification stage utilizes FC layers or global average pooling, followed by a Softmax operator or alternatives like SVM, for predicting class memberships in input patterns. Subsequent sections detail three CNN architectures tailored for HSC: i) Spectral CNN, ii) Spatial CNN, and iii) Spectral-spatial CNN.
6.a Spectral CNN
Spectral CNNs exclusively process 1D spectral data $x_{i}\in\mathcal{R}^{B}$, where $B$ denotes the original or reduced spectral bands (Yu et al.,, 2022). Tailored for HS data, they treat each pixel’s spectral information as a 1D vector, capturing spectral intensity at different wavelengths. Unlike conventional 2D CNNs for image classification, these models consider HSIs as 1D sequences, where each element represents spectral intensity. The input is the pixel’s spectral profile, either the original full-dimensional vector or a dimensionality-reduced representation, presenting a focused approach to handle spectral dimensionality.
@@
A crucial element in spectral CNNs is the convolutional layer, applying filters to capture local spectral patterns and generate feature maps by element-wise operations. Commonly, pooling layers, such as max or average pooling, downsample and reduce dimensionality. Additional layers like fully connected or recurrent layers can be added for enhanced discriminative power. In (Wu et al., 2022a, ), a CNN addressed overfitting by using $1\times 1$ convolutional kernels, increased dropout rates, and a global average pooling layer instead of fully connected layers. For handling high correlation among HSI bands, Gao, et al. (Gao et al.,, 2023) introduced a CNN architecture transforming the 1D spectral vector into a 2D feature matrix with cascading $1\times 1$ and $3\times 3$ convolutional layers. This design allowed feature reuse across different spectral bands. Similarly, Li et al. (Li et al., 2024h, ) employed a global average pooling layer to reduce trainable parameters and extract high-dimensional features.
Wu et al. (Wu and Prasad, 2017a, ) proposed a hybrid model combining convolutional and recurrent layers for extracting both position-invariant middle-level features and spectral-contextual details. Similarly, Jin et al. (Jin et al.,, 2018) proposed a hybrid architecture that effectively classified healthy and diseased wheat heads by transforming spectral information into a 2D structure for improved analysis. In spectral-based identification of rice seed varieties, Qiu et al. (Qiu et al.,, 2018) demonstrated CNN’s superiority over SVM and KNN, showcasing its effectiveness in accurately identifying rice seed varieties based on spectral characteristics. A similar application in (Wu et al.,, 2018) identified various Chrysanthemum varieties using CNN on spectral data represented by the first five Principal Components (PCs) obtained through PCA, a widely-used dimensionality reduction method. In medical HSI, Huang et al. (Huang et al.,, 2018) used PCA as a preprocessing step, fusing CNN kernels with Gabor kernels for classification, highlighting the efficacy of hybrid architectures and the superiority of CNN over traditional algorithms in spectral-based classification tasks for HSC.
Advantages: Firstly, Spectral CNN captures local spectral patterns effectively through convolutional operations, facilitating discriminative feature extraction (Yu et al.,, 2022). Secondly, by integrating pooling operations, spectral CNNs manage high-dimensional HS data, reducing computational complexity. Thirdly, the inclusion of fully connected or recurrent layers enables learning high-level representations and capturing spectral-contextual dependencies, enhancing classification accuracy.